# Generative Physical AI in Vision: A Survey

### v1

@@ Abstract
Generative Artificial Intelligence (AI) has rapidly advanced the field of computer vision by enabling machines to create and interpret visual data with unprecedented sophistication. This transformation builds upon a foundation of generative models to produce realistic images, videos, and 3D/4D content. Traditionally, generative models primarily focus on visual fidelity while often neglecting the physical plausibility of generated content. This gap limits their effectiveness in applications requiring adherence to real-world physical laws, such as robotics, autonomous systems, and scientific simulations. As generative AI evolves to increasingly integrate physical realism and dynamic simulation, its potential to function as a “world simulator" expands—enabling the modeling of interactions governed by physics and bridging the divide between virtual and physical realities. This survey systematically reviews this emerging field of physics-aware generative AI in computer vision, categorizing methods based on how they incorporate physical knowledge—either through explicit simulation or implicit learning. We analyze key paradigms, discuss evaluation protocols, and identify future research directions. By offering a comprehensive overview, this survey aims to help future developments in physically grounded generation for vision. The reviewed papers are summarized at https://github.com/BestJunYu/Awesome-Physics-aware-Generation.
@@ 1 Introduction
Generative learning has long been a foundational pillar of modern computer vision, addressing key challenges in understanding, synthesizing, and manipulating visual data. Over the past decade, this field has witnessed rapid evolution of diverse generative models, including Variational Autoencoders (VAEs)[1, 2], Generative Adversarial Networks (GANs)[3], Diffusion Models (DMs)[4, 5, 6], Neural Radiance Fields (NeRFs)[7], Gaussian Splatting (GS)[8], and Visual Autoregressive Models (VARs)[9]. These models have continually pushed the boundaries of generative learning, leveraging increasingly powerful architectures to capture the underlying distributions of visual data. The goal is to enable machines to reason about the visual world in ways that mirror human creativity and understanding, by imagining novel instances of visual content under unseen scenarios. Among these advancements, diffusion models have emerged as particularly noteworthy for their ability to produce highly realistic outputs. By iteratively refining random noise through learned denoising processes, diffusion models demonstrate exceptional robustness and versatility, making them a cornerstone of recent generative methodologies. The applications of generative models span diverse modalities of visual contents, including image generation with semantics understanding, video generation with dynamic temporal understanding, 3D content generation with enhanced spatial understanding[10, 11, 12], and 4D contents with more complex and combined understanding[13, 14, 15, 16, 17, 18, 19]. These advancements underscore the growing potential of generative learning in the broad landscape of increasingly sophisticated visual tasks.
Figure 1: The picture of generative AI in vision and the role of physics-aware generation on the path towards building a generative world simulator.
@@
Among these different visual modalities, video generation has recently garnered significant attention within the field of generative learning, which serves as a much more challenging testbed for scaling up large generative models to handle higher dimensional data. This complexity stems not only from the spatial intricacies of individual frames but also from the temporal coherence required across sequences. Numerous commercial video generation models have been developed and have captured widespread public interest, such as OpenAI’s Sora[20], Google’s Veo2[21], Tencent’s Hunyuan[22], and Kuaishou’s Kling[23]. Video generation has been intensively studied in a range of formulations and settings, spanning from the vanilla unconditional generation[24, 25], to image-to-video generation[26, 27, 28, 29, 30, 31, 32, 33], text-to-video generation[24, 26, 34, 35, 25, 29, 36, 30, 37, 30], video-to-video generation[38, 39], and video editing or customization[40, 41, 42, 43]. Each of these settings addresses unique challenges, from preserving temporal continuity to incorporating semantic guidance from textual or visual inputs.
@@
More importantly, video occupies a pivotal role in the future of generative AI in vision. The vast amount of video data available online encapsulates a rich repository of information about the real world, making video a medium through which generative AI can learn to model complex real-world phenomena. In this context, video can be regarded as a “language" of real-world decision making, offering the potential to bridge the digital and physical domains[44]. Video generation is posed to provide a unified interface as a “world model”[45] processing physical knowledge, analogous to how textual large language models (LLMs) process abstract knowledge. Such a model could facilitate a large number of downstream tasks, including autonomous driving, scientific simulation, robotics[46, 47, 48, 49, 50], and other forms of embodied intelligence. To realize this potential, the generation process should be capable of interacting with external controls from humans or other systems. This interactivity facilitates dynamic decision-making and the optimization of outcomes based on interactions, giving rise to what can be described as generative interactive environments[51, 44, 52, 53]. Video generation has been incorporated with various interactive control signals, such as motion vectors or trajectories[54, 55, 56, 57, 58], hand masks[59], latent actions[53, 60], robot operations[47], camera motions[61], demonstrations[62], and natural language descriptions[63, 64, 65]. These interactive elements highlight the versatility and adaptability of generative video models, paving the way for their evolution into world models.
@@
However, a critical gap remains in the transition from generation to robust world modeling: the ability to faithfully understand and replicate real-world physics[66] (Fig. 1). Current state-of-the-art models are primarily optimized for visual realism in the pixel space, rather than the physical plausibility in the entity or concept space. For generative models to serve as simulators of the physical world, they must incorporate a deep understanding of physical laws, such as dynamics, causality, and material properties. This physics awareness is essential for going beyond generating visually appealing outputs to ensure that the content also aligns with the constraints and behaviors of the physical world.
Therefore, we provide this survey as a timely and comprehensive review of existing literature that attempts to embed physics-awareness into generative models. By examining these efforts, we hope to highlight the progress made so far, provide a clear structure of paradigms, and identify potential future directions.
@@
Survey Scope. The scope of this survey is about generative models in computer vision that enhance the physical awareness of the generation outputs. Therefore, we exclude from our scope the literature that incorporates physical principles as prior knowledge or inductive biases in model or neural architecture design, such as Physics-Informed Neural Networks (PINNs)[67, 68], even if the task is related to generative learning, e.g., [69, 70, 71]. We focus on generative tasks, and thus exclude image processing tasks such as de-blurring, de-hazing, and enhancement from our scope, though we notice that a large body of these works incorporate physics. To concentrate on computer vision, pure graphic and rendering research combined with physical simulation is also excluded from our scope.
@@
Comparison to Other Surveys. As delineated in our scope, this survey is orthogonal to existing surveys on physics-informed machine learning[72], physics-informed computer vision[73], and physics-informed artificial intelligence[74], since they emphasize the model design aspect with physical prior knowledge. Our survey focuses on generation with physical awareness, thus being distinct from existing surveys on generative models[75], diffusion models[76, 77], video diffusion models[78], diffusion-based video editing[79] in a general sense. Our survey also has a different scope compared to those in specific domains, such as human video or motion generation[80, 81, 82].
@@ 2 Formulation
In this section, we first provide a definition of physics-aware generation and related concepts, such as physical simulation and physical understanding. Based on the definition, we further identify the common paradigms of incorporating physics into generative models in vision, to provide a structural perspective for later sections of this survey.
@@ 2.1 Definitions
Let $P_{\theta}$ denotes a physical simulation model with physical parameters $\theta$, and $G$ represents a generative model, we can then provide the following definitions:
@@ 2.2 Formulation of Physics-Aware Generation
Based on the above concepts, we further define the concept of physics-aware generation (PAG) as the generation process with a strong understanding of real-world physics. We divide physics-aware generation into two major categories, i.e., physics-aware generation with explicit physical simulation (PAG-E), and physics-aware generation without explicit physical simulation (PAG-I). This division is based on whether the generative model utilizes a physical simulation model explicitly to improve physical awareness. Table I provides a comparison of these concepts.
@@
Simulation in Generation (SiG): In this paradigm, the simulation model is incorporated as a part or a sub-module of the generation model:
@@
Generation and Simulation (GnS): In this paradigm, the generation and the simulation happen simultaneously, such as by using a shared model $M$:
@@
Simulation-Constrained Generation (ScG): In this paradigm, the simulation model is used to provide constraints or knowledge to the generation model:
$G(X)\rightarrow X^{\prime}\;\;subject\;to\;\;P_{\theta}(X)\rightarrow X^{\prime}.$ (7)
Generation-Constrained Simulation (GcS): In this paradigm, the generation model is used to provide constraints or knowledge for the simulation model:
$P_{\theta}(X)\rightarrow X^{\prime}\;\;subject\;to\;\;G(X)\rightarrow X^{\prime}.$ (8)
Simulation-evaluated Generation (SeG): In this paradigm, physical simulation is used to evaluate the generative model or the generated content is intended for deployment in simulation environments.
These paradigms are illustrated in Fig. 2. In the following sections, we will cover physical simulation in Section 3, physical understanding in Section 4, physical-aware generation in Section 5, and physical evaluation in Section 6.
Figure 2: Paradigms of incorporating explicit physical simulation with generative models.
Figure 3: Examples of physical materials, simulation methods, and physical engines used in physics-aware generation. Images of physical materials are generated by GPT-4o.
@@ 3 Physical Simulation
Physical simulation refers to the use of simulated environments or models that mimic real-world physical systems, allowing the generative models to learn and infer without direct interaction with the physical world. In the following, we summarize the core elements of physical simulations commonly used in physics-aware generation research, along the dimensions of physical material properties, simulation methods, and off-the-shelf physical engines.
3.1 Physical Materials
In this section, we introduce several physical materials widely used in physical simulation when integrated with physics-aware generation, as detailed below.
Rigid Body. This refers to a material or object that does not deform when subjected to applied forces [83, 84, 85, 86, 87, 88, 89, 90, 91]. In physics-aware generative models, rigid bodies are often used to represent components that maintain their shape and volume. These models simulate rigid body dynamics, including translation, rotation, and collision, to predict the movement of objects.
Soft Body. Soft body is a material [92, 85] that can undergo large deformations when subjected to external forces such as biological tissues, rubber, gels, cloth, and polymers. They exhibit properties like elasticity, viscoelasticity, and non-linear behavior, and they often have complex internal structures.
Newtonian Fluids. Newtonian fluids [93, 89, 94, 95] are fluids whose viscosity (resistance to flow) remains constant regardless of the applied stress or the rate of flow. Common examples include water, air, and most gases. In the modeling process, fluid motion can be described using standard equations like the Navier-Stokes equations.
Non-Newtonian Fluids. These are fluids whose viscosity changes when a stress is applied or when the rate of shear changes [93, 96, 94]. Examples include ketchup, cornstarch in water, and blood. Non-Newtonian fluids can exhibit behaviors such as shear-thinning (viscosity decreases with increasing shear rate), shear-thickening (viscosity increases with increasing shear rate), or even more complex behaviors like viscoelasticity.
Plasticine. Plasticine [93, 88, 86, 97, 98, 94, 91], as a soft and viscoelastic material, is often modeled using viscoelastic and soft body physics approaches. These methods capture both the elastic (rebound) and viscous (flowing or slow deformation) aspects of plasticine’s behavior, making it suitable for simulations of materials that deform and retain new shapes.
Elastic Entities. Elastic materials [94, 88, 98, 99, 93, 84, 100, 86, 96, 97, 101, 91] are characterized by their ability to return to their original shape after deformation, following a linear or nonlinear stress-strain relationship.
Granular Media. Granular media [91, 96], such as sand [94, 102, 103, 93] and snow, behave in complex ways that combine both solid and fluid properties. Their behavior is governed by discrete particle interactions, which are often modeled using Discrete Element Method (DEM) or adapted continuum models that capture phenomena like flow, compaction, and jamming.
Metals. Metals [96, 104, 102] have the ability to undergo elastic and plastic deformations. They are described using elasticity and plasticity models, including work-hardening and yield criteria. These models are crucial for simulating metal behavior under different loads.
Solid. Solids [87, 88, 89] are essential components in physical models due to their definite shape, strength, and mechanical properties. Whether it is modeling stress and strain, plasticity, or thermal expansion, understanding the behavior of solid materials is crucial for accurate simulations and predictions.
Cloth Materials. Cloth [90, 89] exhibits unique behavior due to its flexibility, anisotropy, and resistance to deformation. In simulations, it is usually modeled using techniques like mass-spring systems or particle-based systems to simulate deformations and interactions with forces like wind and gravity.
Articulated Bodies. Articulated bodies [105, 104, 95] are central to many physical models, especially in systems with connected, movable parts like robots, human bodies, and mechanical systems. Their kinematic and dynamic behavior is complex, accounting for motion, forces, collisions, and constraints. These models are used in fields where understanding the interaction of interconnected bodies is crucial.
Thin-shell Objects. Thin-shell objects [106, 105] in physical models refer to structures that have a small thickness compared to their other dimensions (length and width) but still exhibit significant mechanical behavior that needs to be modeled accurately.
3.2 Physical Parameters
Physical materials are substances that have specific properties, such as density, elasticity, and thermal conductivity, which define their behavior under different conditions [101, 107, 97, 108, 95, 83, 109, 110, 99, 94, 111, 86, 105, 92, 93, 88, 89, 87, 102, 98, 104]. These properties are quantified by physical parameters, which are measurable values that describe the material’s response to external forces, heat, or other stimuli. For example, the Young’s modulus of a material defines its stiffness, while thermal conductivity measures its ability to conduct heat. The relationship between materials and their physical parameters is crucial for predicting how materials will behave. The definition of physical parameters is usually correlated with the selection of the physical material and the simulation method. We provide some examples of common parameters for each physical material as shown in Figure 3.
3.3 Simulation Methods
Below we introduce methods for simulating and analyzing physical dynamics, which appear in the literature of physics-aware generation reviewed in this survey.
Continuum Mechanics-based Method (CMBM). Continuum mechanics-based method [106] solves the governing partial differential equations that describe the behavior of materials as continuous media, focusing on stress, strain, and motion in solids, fluids, and gases. These simulations use numerical methods like FEM or Computational Fluid Dynamics to approximate solutions for complex, real-world systems involving large deformations, fluid flow, and heat transfer.
Material Point Method (MPM). The MPM [103, 91, 96, 86, 88, 97, 84, 101, 98, 93, 94, 107, 112] is a hybrid simulation technique that combines Lagrangian particles with an Eulerian grid to model the behavior of materials. In this method, particles carry material properties such as mass, velocity, and stress, while a background grid is used to solve the governing equations of motion. The particles interact with the grid, allowing MPM to efficiently handle large deformations, phase transitions, and multiphase flow while capturing both material behavior and complex interactions.
Finite Element Method (FEM). FEM [110, 96, 103, 100] discretizes a physical domain into small, interconnected elements and solves the governing equations of motion or energy using approximation functions within each element. These element-level equations are then assembled into a global system, which is solved to obtain the behavior of the entire system under various loads and boundary conditions.
Position-Based Dynamics (PBD). PBD [91, 102] is a physics simulation method that focuses on solving constraints directly in terms of particle positions, rather than velocities or forces. The technique iteratively adjusts particle positions to satisfy physical constraints, such as distance, volume preservation, or collision, ensuring stability and efficiency, especially for real-time applications involving deformable objects and soft materials.
Eulerian Method (EM). EM [93, 112, 94, 95] is a grid-based method where the physical domain is fixed, and quantities like velocity, pressure, or temperature are computed at fixed points in space. As the material flows through the grid, EM solves the governing equations (e.g., conservation of mass, momentum) at each grid point to model the evolution of physical fields over time.
Lagrangian Method (LM). LM [93, 112, 100, 109, 94, 95] follows individual particles or material points as they move through space and time, solving the equations of motion based on their positions, velocities, and forces. In practice simulation, LM is particularly useful for tracking the deformation and motion of materials with moving boundaries or interfaces, such as in fluid dynamics or solid mechanics.
3D Spring-Mass Model. This is a physical simulation technique where a deformable object is discretized into masses connected by springs [99], each spring representing elastic forces that govern the object’s deformation. The motion of the masses is computed over time by solving the equations of motion using numerical methods, accounting for forces such as spring tension, damping, and external interactions.
Newtonian Dynamics. The Newtonian Dynamics represents a physical simulation method [103] that models the motion of objects based on Newton’s laws of motion, where the force acting on an object is equal to its mass times its acceleration. It uses numerical integration techniques to solve the equations of motion over time, simulating the object’s response to forces such as gravity, collisions, and applied external forces.
$\Phi$-Flow. This is a physics-based simulation method that uses a level-set function $\Phi$ to represent fluid interfaces and track free-surface flows or multiphase interactions [95]. The method solves the Navier-Stokes equations for fluid dynamics while evolving the level-set function to handle complex interfaces and simulate realistic fluid behaviors.
3.4 Physics Engines and Platforms
To perform physics simulations, several popular engines and platforms have been developed, as summarized below.
Bullet Physics. This is an open-source physics engine widely used in games, animation, and robotics simulations [113]. It enables to support rigid body dynamics, soft body dynamics, collision detection, and physical constraints.
Havok Physics. Havok [114] is a commercial physics engine widely used in games and real-time simulations. It offers efficient rigid body, fluid, cloth, and vehicle physics simulations.
NVIDIA PhysX. This is a physics engine developed by NVIDIA that supports GPU-accelerated real-time simulation [115]. It includes particle systems, cloth, rigid body dynamics, and fluid simulations, optimized with high-performance.
Unity Physics. The built-in physics engine of Unity [91, 116] supports both 2D and 3D physics simulations. It provides collision detection, rigid body dynamics, cloth, and soft body physics.
Unreal Engine Physics (Chaos Physics). The physics engine in Unreal [90] supports large-scale destruction and fluid/cloth simulations, also providing realistic rigid body and soft body physics.
Open Dynamics Engine. This an open-source physics engine [117] focusing on rigid body dynamics and collision detection, widely used in games, robotics, and physical simulation systems.
Box2D. Box2D [118] is a lightweight, open-source 2D physics engine designed specifically for 2D simulation, offering efficient rigid body dynamics and collision detection.
Mantaflow. Mantaflow [119] is an open-source fluid and smoke simulation engine, with the capability to support high-quality fluid, fire, and gas simulations.
Simulink/Matlab (Simscape Multibody). This is a platform within Matlab used for multi-body dynamics and mechanical system simulations [120]. It is highly suitable for detailed modeling and simulation of mechanical systems, robotics, and vehicles.
Blender. This is a free and open-source 3D creation suite [103, 89] that supports modeling, sculpting, animation, rendering, and simulation. It offers a powerful and versatile toolset for productions and scientific simulations.
Isaac Gym. Isaac Gym [109] is a high-performance physics simulation platform developed by NVIDIA, designed specifically for training and testing robotic systems using reinforcement learning. It provides a GPU-accelerated framework that allows to simulate multiple environments in parallel, enabling faster training of agents.
Vortex Studio. This is a high-performance simulation platform primarily designed for robotics, vehicles, and mechanical systems [121]. Its most attractive feature is its ability to offer real-time physics simulation, with a focus on large systems and complex dynamics.
PyBullet. PyBullet is a Python interface for Bullet Physics [122, 123, 90, 108], commonly used in robotics and reinforcement learning, providing simulations of rigid body dynamics in machine learning environments.
Gazebo. Gazebo is a robot simulation platform that provides high-fidelity physics and sensor modeling [124]. It integrates with Robot Operating System for realistic robotics simulations.
Genesis. Genesis [125] is an open-source platform for embodied artificial intelligence that provides a high-fidelity physics environment to simulate robots and their interactions with the physical world. It focuses on helping agents learn through physical interaction with environments, supporting tasks like manipulation, locomotion, and real-world decision-making.
Taichi. Taichi [103] is a platform combining the Taichi programming framework to enhance the efficiency and accuracy of physical system simulations. By integrating optimization and learning algorithms, it enables adaptive modeling, real-time simulation adjustments, and data-driven insights for robotics, virtual environments, and scientific simulation.

### v2

@@ Abstract
Generative Artificial Intelligence (AI) has rapidly advanced the field of computer vision by enabling machines to create and interpret visual data with unprecedented sophistication. This transformation builds upon a foundation of generative models to produce realistic images, videos, and 3D/4D content. Conventional generative models primarily focus on visual fidelity while often neglecting the physical plausibility of the generated content. This gap limits their effectiveness in applications that require adherence to real-world physical laws, such as robotics, autonomous systems, and scientific simulations. As generative models evolve to increasingly integrate physical realism and dynamic simulation, their potential to function as “world simulators" expands. Therefore, the field of physics-aware generation in computer vision is rapidly growing, calling for a comprehensive survey to provide a structured analysis of current efforts. To serve this purpose, the survey presents a systematic review, categorizing methods based on how they incorporate physical knowledge—either through explicit simulation or implicit learning. It also analyzes key paradigms, discusses evaluation protocols, and identifies future research directions. By offering a comprehensive overview, this survey aims to help future developments in physically grounded generation for computer vision. The reviewed papers are summarized at https://tinyurl.com/Physics-Aware-Generation.
@@ 1 Introduction
Generative learning has long been a foundational pillar of computer vision, addressing key challenges in understanding, synthesizing, and manipulating visual data. Over the past decade, this field has witnessed rapid evolution of diverse generative models, including Variational Autoencoders (VAEs)[1, 2], Generative Adversarial Networks (GANs)[3], Diffusion Models (DMs)[4, 5, 6], Neural Radiance Fields (NeRFs)[7], Gaussian Splatting (GS)[8], and Visual Autoregressive Models (VARs)[9]. These models have consistently pushed the boundaries of generative learning, leveraging increasingly powerful architectures to capture the underlying distributions of visual data. The goal is to enable machines to reason about the visual world in ways that mirror human creativity and understanding, by imagining novel instances of visual content under unseen scenarios. Among these advancements, diffusion models have emerged particularly noteworthy for their ability to produce high-fidelity outputs. By iteratively refining random noise through learned denoising processes, diffusion models demonstrate exceptional robustness and versatility, making them a cornerstone of recent generative methodologies. The applications of generative models span diverse modalities of visual contents, including image generation with semantics understanding, video generation with dynamic temporal understanding, 3D content generation with enhanced spatial understanding[10, 11, 12], and 4D contents with more complex spatiotemporal understanding[13, 14, 15, 16, 17, 18, 19]. These advancements underscore the growing potential of generative learning in the broad landscape of increasingly sophisticated visual tasks.
Figure 1: Overview of the progression of generative AI in vision and the role of physics-aware generation in building a world simulator. The figure shows the evolution from conventional focuses in generative tasks on semantic, temporal, and spatial understanding to incorporating interactivity and physical realism. Physics-aware generation enables models to simulate real-world dynamics, advancing toward general-purpose world models for various applications.
Figure 2: Examples of generated videos with poor and good physical awareness. Examples are from WISA [20] and Zhao et al [21].
@@
Among these visual modalities, video generation has recently gained significant attention in generative learning, presenting a far more challenging testbed for scaling up large generative models to handle higher dimensional data. This complexity stems not only from the spatial intricacies within individual frames but also from the temporal coherence required across multiple frames. Numerous video generation models have been developed and have captured widespread public interest, such as OpenAI’s Sora[22], Google’s Veo2[23], Tencent’s Hunyuan[24], Kuaishou’s Kling[25], and Nvidia’s Cosmos [26]. Video generation has been intensively studied in a range of formulations and settings, spanning from the vanilla unconditional generation[27, 28], to image-to-video generation[29, 30, 31, 32, 33, 34, 35, 36], text-to-video generation[27, 29, 37, 38, 28, 32, 39, 40, 33], video-to-video generation[41, 42], and video editing or customization[43, 44, 45, 46]. Each of these settings addresses unique challenges, from preserving temporal coherence to incorporating semantic guidance from textual or visual inputs.
@@
Video has a pivotal role in generative AI in computer vision, since the vast amount of video data available online encapsulates a rich repository of information about the real world, making video a medium through which generative AI can learn to model complex real-world phenomena. In this context, video can be regarded as an implicit physical model of the world, offering the potential to bridge the digital and physical domains[47], and paving the way for real-world decision-making. Such a model could facilitate a large number of downstream tasks, including autonomous driving, scientific simulation, robotics[48, 49, 50, 51, 52], and other forms of embodied intelligence. To realize this potential, the generation process must be capable of interacting with external controls from humans or other systems. This interactivity facilitates dynamic decision-making and the optimization of outcomes based on interactions, giving rise to what can be described as generative interactive environments[53, 47, 54, 55]. Video generation has been incorporated with various interactive control signals, such as motion vectors or trajectories[56, 57, 58, 59, 60], hand masks[61], latent actions[55, 62], robot operations[49], camera motions[63], demonstrations[64], and natural language descriptions[65, 66, 67]. These interactive elements highlight the versatility and adaptability of generative video models, showing the potential to evolve into world models.
@@
However, a critical gap remains in the transition from generation to robust world modeling, i.e., the ability to faithfully understand and replicate real-world physics[68] (Fig. 1). Current state-of-the-art models are primarily optimized for visual realism in the pixel space, rather than physical plausibility in the entity or concept space (Fig. 2). For generative models to serve as physical world simulators, they must incorporate a deep understanding of physical laws, such as dynamics, causality, and material properties. This physics awareness is essential for going beyond generating visually appealing outputs to ensuring that the content also aligns with the constraints of the physical world. The field of physics-aware generation is rapidly expanding, yet it remains fragmented, encompassing a wide range of diverse tasks, lacking clear definitions and standardized evaluation protocols. Therefore, we provide this survey as a timely and comprehensive review of existing literature to streamline research efforts. By examining existing efforts, we aim to highlight the progress made so far, provide a structure of paradigms, and identify potential future directions.
@@
Survey Scope. This is the first survey with a scope of generative models in computer vision that enhance the physical awareness of the generation outputs. We shift away from discussing methods that incorporate physical principles as prior knowledge or inductive biases in the model or neural architecture design, such as Physics-Informed Neural Networks (PINNs)[69, 70], even if the task is related to generative learning, e.g., [71, 72, 73]. Moreover, we focus on generative tasks, and thus exclude image processing tasks such as de-blurring, de-hazing, and enhancement from our scope, though we notice that a large body of these works incorporate physics. Finally, to keep the survey fully focused on computer vision, pure graphics research involving physical simulation is also excluded.
@@
Comparison to Other Surveys. This survey is orthogonal to existing surveys on physics-informed machine learning[74], physics-informed computer vision[75], and physics-informed artificial intelligence[76], since they emphasize the model design aspect with physical prior knowledge rather than enhancing the physical realism of generation outputs. Our survey focuses on physics-aware generation, thus being distinct from existing surveys with general scopes on generative models[77], diffusion models[78, 79], video diffusion models[80], diffusion-based video editing[81]. Our survey also has a different scope compared to those in specific domains, such as human video or motion generation[82, 83, 84]. A recent, subsequent survey [85] has a similar scope, but from a different perspective of cognitive evolution processes of physics in video generation models.
@@ 2 Formulation
We first provide definitions of physics-aware generation and related concepts, such as physical simulation and physical understanding. Based on these definitions, we further identify the common paradigms of incorporating physics into generative models in computer vision, to provide a structural perspective for later sections of this survey.
@@ 2.1 Definitions
Let $P_{\theta}$ denote a physical simulation model with physical parameters $\theta$, and $G$ represent a generative model, we can then provide the following definitions:
@@ 2.2 Formulation of Physics-Aware Generation
Based on the above concepts, we further define the concept of physics-aware generation (PAG) as the generation process with a strong understanding of real-world physics. We divide physics-aware generation into two major categories, i.e., physics-aware generation with explicit physical simulation (PAG-E), and physics-aware generation with implicit learning (PAG-I). This division is based on whether the generative model utilizes a physical simulation model explicitly to improve physical awareness. Table I provides a comparison of these concepts.
@@
Figure 3: Illustration of six paradigms for integrating explicit physical simulation into generative models, highlighting structural differences in terms of the interaction between the generative process $G$ and the physics simulation model $P$.
@@
Simulation in Generation (SiG): The simulation model is incorporated as a part or a sub-module of the generation model:
@@
Generation and Simulation (GnS): The generation and the simulation happen simultaneously, such as by using a shared model $M$:
@@
Simulation-Constrained Generation (ScG): The simulation model is used to provide constraints or knowledge to the generation model:
$G(X)\rightarrow X^{\prime}\;\;\mathrm{subject\;to}\;\;P_{\theta}(X)\rightarrow X^{\prime}.$ (7)
Generation-Constrained Simulation (GcS): The generation model is used to provide constraints or knowledge to the simulation model:
$P_{\theta}(X)\rightarrow X^{\prime}\;\;\mathrm{subject\;to}\;\;G(X)\rightarrow X^{\prime}.$ (8)
Simulation-evaluated Generation (SeG): The physical simulation is used to evaluate the generative model or the generated content is intended for deployment in simulation environments.
These paradigms are illustrated in Fig. 3. We cover physical simulation in Section 3, physical understanding in Section 4, physical-aware generation in Section 5, and physical evaluation in Section 6.
Figure 4: Overview of key components in physical simulation relevant to physics-aware generation, i.e., physical materials, simulation methods, and physical engines. Physical materials define the types of entities being modeled with different assumptions and constraints, with illustrations generated by GPT-4o. Simulation methods provide the computational tools to model the dynamics of these materials under physical laws. Physics engines serve as practical and off-the-shelf platforms implementing these simulations. Examples are given for each component for illustration.
@@ 3 Physical Simulation
Physical simulation refers to the use of simulated environments or models that mimic real-world physical systems, allowing the generative models to learn and infer without direct interaction with the physical world. In Fig. 4, we summarize the core elements of physical simulations commonly used in physics-aware generation research, along the dimensions of physical material properties, simulation methods, and off-the-shelf physical engines. Detailed descriptions are given in the Appendix.